The perimeter of a rectangle is 36 yards. The width is 18 yards less than twice the length. FInd the length and width of the rectangle.
Answer by checkley77(12844)

P=2L+2W
W=2L-18
36=2L+2(2L-18)
36=2L+4L-36
6L=36+36
6L=72
L=72/6
L=12 ANS.
W=2*12-18
W=24-18=6 ANS FOR THE WIDTH.
PROOF:
36=2*12+2*6
36=24+12
36=36